A and b stand in a ring with 9 persons.if the arrangement of 11 persons is at random then the probability that there are exactly 4 persons between a and b is

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
Fix the position of A, then B can be sit in 10 positions, so total possible outcome = 10
Now, exactly 4 persons are sitting between them. This can be done in two ways as shown in figure, so favourable outcomes = 2
So, probability = 2/10 = 1/5
 
 
Regards
Arun (askIITians forum expert)